Find the equation of the cylinder whose
axis is x = y = - z and radius is 2.
1
Expert's answer
2020-12-28T19:11:19-0500

Let "O(0,0,0), A(1,1,-1)" be two points on the cylinder axis, "X(x,y,z)" be any point on the cylinder surface, "P(p_x, p_y, p_z)" be the orthogonal projection of the point X onto the axis. Then the length |XP| is a distance from X to the axis.

The vector "OP" is a projection of "OX" , it is equal to "OP = OA<OX,OA>\/|OA|^2" .

"<OX,OA> = x + y - z"

"|OA|^2 = 1^2 + 1^ 2 + (-1)^2 = 3"

"PX =OX - OP = (x-p_x,y-p_y,z-p_z)"

"x-p_x = x - (x+y-z)\/3 = (2x-y+z)\/3"

"y-p_y = y - (x+y-z)\/3 = (2y-x+z)\/3"

"z-p_z = z + (x+y-z)\/3 = (x+y+2z)\/3"

"4 =|PX|^2=((2x-y+z)^2 + (2y-x+z)^2+(x+y+2z)^2)\/9"

"36 = 6(x^2+y^2+z^2-xy+yz+zx)"

"x^2+y^2+z^2-xy+yz+zx=6"

This is the equation of the cylinder with the axis "OA" and radius 2.